[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Fri Jun 17, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Fri Jun 17 17:57:08 UTC 2016


HSU Shift Log for Fri Jun 17 2016 13:57:03 GMT-0400 (EDT)

Shift: H

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues: None

N-AWIPS/AWIPS2 issues: I modified the satellite SPFs for work spaces 2 and  
3 to set the FILTER parameter for the ship/buoy obs to 0.01 to prevent the  
over plotting issue reported earlier. The double-plotting of the surface  
obs in the ABC islands (and elsewhere) is due to loading both the synoptic  
obs and hourly METAR obs at the same time. If you turn one of those sources  
off you can see the other.

Other Techincal issues: None
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160617/a02350f4/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list